# Retrieval-Augmented Generation for Natural Language Processing: A Survey

### v2

@@ 6.2. RAG with Datastore Update
As shown in Figure 4 (d), this scenario involves two stages: updating the knowledge database, then training the retriever and the generator. There are three cases for updating the knowledge database, i.e., updating with trainable embeddings, updating with new values, and updating with new corpus. In the first case, values generally are trainable embeddings and are simultaneously/asynchronously updated with parameters in the RAG (Chen et al., 2022b). The last two cases usually refer to updating the knowledge database with up-to-date information. Taking question-answer corpus as an example, updating with new values refers to updating the answer to existing questions, while updating with new corpus refers to adding new question-answer pairs. To update the value of existing keys requires first querying the existing key-value pairs and then performing in-place updates. For a new corpus, the datastore first needs to perform insertion operations, then rebuilds or updates the indexes for new keys. After updating the datastore, training the retriever and the generator is similar to RAG without datastore update. However, this training step is not always a necessary step, benefiting to the in-context learning capability of LLMs.
@@ 7.4. Question Answering
RAG techniques combine information retrieval with model-based generation, which is highly suitable for QA systems. In particular, open-domain QA systems usually first require searching for knowledge from the Internet or large-scale databases, then generate the corresponding answers according to the retrieved knowledge. Naturally, given similar questions and corresponding answers as demonstrations which are concatenated into inputs (Wang et al., 2022; Li et al., 2023b; Huang et al., 2023c), generators in RAG can learn the pattern between questions and answers and infer what answers should be. For some specific QA tasks where a set of reference documents is given, retrievers in RAG would retrieve the relevant documents for concatenation, and then generators in RAG would read the context then generate the final answers via the self-attention mechanism (Guu et al., 2020; Lee et al., 2023; Ram et al., 2023b; Asai et al., 2023), which is similar to solving a reading comprehension problem. Besides, Fabbri et al. (Fabbri et al., 2020) focus on designing effective templates for re-organizing the concatenated contexts. Baek et al. (Baek et al., 2023) leverage the knowledge graph to retrieve the related facts for the input questions, then feed their concatenation and inputs into the generators. Instead of directly concatenating texts, another branch of works focuses on joining the retrieval embeddings with input embeddings for the encoder-decoder models (Izacard and Grave, 2021; Sachan et al., 2021; de Jong et al., 2023b; Izacard et al., 2023).
@@
Some works incorporate the external knowledge in the hidden states or the final logits of generators. For the fusion in the hidden states, the key is what kind of knowledge should be injected, such as entities (Févry et al., 2020; de Jong et al., 2022), chunks (Borgeaud et al., 2022; Wang et al., 2023e), documents (Cheng et al., 2023a). For the fusion in the logits, most works combine the logits of retrievals and inputs by ensemble techniques (Shi et al., 2023; Guu et al., 2020; Lewis et al., 2020; Mueller et al., 2023).
@@
Instead of designing different knowledge fusions for QA systems, existing works also improve QA systems with RAG from other aspects. Some works (Guo et al., 2019; Paranjape et al., 2022; Lin et al., 2022) use retrieved question-answering pairs as extra training data. Some works optimize the retriever module, e.g., improving the keys’ representation when building the retriever database (Ram et al., 2023a), replacing the indexing with a pre-trained ranking model (Yu et al., 2023b), or enabling retrieving phrases with two queries (Min et al., 2023). Other works focus on accelerating the generation efficiency of RAG. Jong et al. (de Jong et al., 2023a) propose the layer-sparse cross-attention to speed up the decoding. Some works (Asai et al., 2023; Jiang et al., 2023b; Wang et al., 2023c) observe that the retrievals may not always provide useful information during the generation process and learn to determine when to retrieve. Moreover, Sun et al. (Sun et al., 2023) combine the RAG with agents to iteratively reason the final results.
@@ 7.7. Dialogue Systems
Existing works improve the dialogue system with RAG mostly via the concatenation-based methods. Some works (Li et al., 2021; King and Flanigan, 2023; Cheng et al., 2023b) concatenate the retrieved history conversations with current inputs. Other works (Fan et al., 2021; Liu et al., 2023a; Cheng et al., 2023b) first leverage an encoder to encode the history responses, then feed the concatenated embeddings into a decoder to generate new responses.
@@ 8.1. LLM-based Autonomous Agents
LLM-based autonomous agents are intelligent software systems which leverages the power of LLMs to perform tasks without the need for continuous human intervention (Li et al., 2024; Xi et al., 2023; Wang et al., 2024). These agents use LLMs as a brain or controller (Huang et al., 2024), and extend their abilities through multimodal perception (Xie et al., 2024), tool utilization (Schick et al., 2023) and external memory (Packer et al., 2023). Especially, external long-term memory for agents functions as the knowledge datastore in RAG, which provides agents with the capability to incorporate external knowledge over extended periods. Therefore, applying RAG would be benefit to access a broader range of information, improving agents’ decision-making and problem-solving abilities (Zhang et al., 2024b). This section explores how LLM-based agents can leverage RAG from two perspectives.
@@ 8.2. Frameworks
Frameworks like Langchain (LangChain, 2023) and LLaMAindex (Liu, 2022) pose significant impact on enhancing the practical implementation of RAG. Langchain and LLaMAindex exemplify the integration of sophisticated retrieval mechanisms with generative models, facilitating the seamless incorporation of external data into the language generation process. This section will introduce these two representative RAG frameworks in details.
Langchain is a framework designed to augment the capabilities of language models by integrating them with external knowledge sources and databases. It acts as a middleware that facilitates the interaction between language models and various data retrieval systems, enabling more informed and accurate generation of responses. The core functionality of Langchain involves orchestrating the flow of information from external databases into the generative process of language models, enhancing their ability to leverage context and specific knowledge in their responses. This integration plays a crucial role in enabling language models to perform tasks that require access to up-to-date or detailed information that is not contained within the model’s initial training data.
@@ 9. Discussion and Future Direction
Despite the success of the RAG for natural language processing, there are some challenges that should be considered. This paper highlights these challenges to inspire future research and provides possible future research directions in RAG for NLP.
@@ 9.1. Retrieval Quality
The retrieval quality refers to improving the relevance of the information retrieved in RAG, which involves the following four key factors to be designed. The first consideration is determining the optimal key to use in the vector database. This process typically involves subjective decision-making and requires human effort to design effectively. The naive idea is to choose inputs for the given tasks, treating each task as a QA problem.
The second is the choice of embedding model. After determining the key, the next step is leveraging embedding models to convert text into vector representations. Models such as BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), or domain-specific embeddings can be crucial to determine how well nuances and contextual meanings are captured. Adapting the embedding model to better suit specific types of data or queries can significantly enhance retrieval quality. This requires training the model on domain-specific corpora that include the types of queries and documents the system will encounter.
@@ 9.4. RAG Training
As introduced in Section 6, RAG training includes two branch of works, RAG with/without datastore update. For RAG without datastore update, the main challenge is how to jointly optimize all parameters in RAG. This may involves new loss functions with multiple objectives, new optimizations for efficient tuning parameters in retriever and generator, or other training strategies.
For RAG with datastore update, one challenge is how to align the retrieval representations with the generator’s representations. Although the time cost of the update operation in datastore cannot be ignored, some works (Chen et al., 2022b) reduce the update frequency by asychronously updating, thus achieving the alignment of knowledge representation and model’s representation. Another challenge is when to retrain/fine-tune the generator in RAG when new corpus is added. Due to the in-context learning capability of exisitng LLM-based generators and high training overhead, retraining/fine-tuning the generator or directly inferring the generator becomes a challenging choice for different scenarios. Recently, some efficient training strategies (Hu et al., 2022; Dettmers et al., 2023) have been proposed to accelerate the fine-tuning process, which can be taken into considerations.
@@ 10. Conclusion
In this survey, we delve into the development of RAG within the field of natural language processing. First, this paper introduces the components of RAG and their functionalities. Subsequently, this paper elaborates on each step involved in retriever, discussing the diverse techniques. Furthermore, this paper categorizes the retrieval fusions, evaluating the strengths and weaknesses inherent of each retrieval fusion techniques. Besides, this paper discusses the RAG training, including RAG with/without datastore update. Then, this paper explores how RAG can be adapted for various NLP tasks and provides practical applications of RAG in real-world scenarios. Conclusively, this paper identifies ongoing challenges and suggests directions for future research to foster advancements in this evolving area.

### v3

@@ 6.2. RAG with Datastore Update
As shown in Figure 4 (d), the scenario involves two stages: updating the knowledge database, then training the retriever and the generator. There are three cases for updating the knowledge database, i.e., updating with trainable embeddings, updating with new values, and updating with new corpus. In the first case, values generally are trainable embeddings and are simultaneously/asynchronously updated with parameters in the RAG (Chen et al., 2022b). The last two cases usually refer to updating the knowledge database with up-to-date information. Taking question-answer corpus as an example, updating with new values refers to updating the answer to existing questions, while updating with new corpus refers to adding new question-answer pairs. To update the value of existing keys requires first querying the existing key-value pairs and then performing in-place updates. For a new corpus, the datastore first needs to perform insertion operations, then rebuilds or updates the indexes as new keys are added. After updating the datastore, training the retriever and the generator is similar to RAG without datastore updates. However, this training step is not always necessary, thanks to the in-context learning capability of LLMs.
7. RAG Evaluation and Benchmark
Retrieval-Augmented Generation (RAG) systems integrate the capabilities of large language models (LLMs) with external knowledge retrieval to advance text generation tasks in natural language processing (NLP). Evaluating RAG systems requires a dual focus on retrieval quality and generation performance to ensure the system effectively leverages external knowledge and produces accurate, contextually appropriate outputs. Recent research has introduced domain-specific benchmarks and comprehensive evaluation frameworks to address these challenges.
For retrieval quality, Pipitone et al. (Pipitone and Alami, 2024) proposed a benchmark tailored for RAG systems in the legal domain, employing key retrieval metrics such as Precision@K and Recall@K to assess the relevance and ranking of retrieved documents. Adlakha et al. (Adlakha et al., 2024) expanded the evaluation scope by incorporating correctness, faithfulness, and human evaluation to assess retrieved information, alongside novel token-overlap metrics for finer-grained analysis.
On the generation side, Hui et al. (Hui et al., 2024) introduced a benchmark suite for RAG systems in real-world document analysis, utilizing metrics such as accuracy, F1 score, and exact match score to evaluate answer quality. Similarly, Xiong et al. (Xiong et al., 2024) developed a medical RAG benchmark to systematically compare the performance of medical RAG systems, emphasizing domain-specific evaluation criteria.
Recent advancements have further enriched the evaluation landscape by proposing more systematic and comprehensive methodologies. Chen et al. (Chen et al., 2024) identified four fundamental RAG abilities—noise robustness, negative rejection, information integration, and counterfactual robustness—as key evaluation dimensions. Saad-Falcon et al. (Saad-Falcon et al., 2024) introduced ARES, an automated RAG evaluation system, which assesses RAG systems using context relevance, answer faithfulness, and answer relevance. Building on this, Friel et al. (Friel et al., 2024) extended the evaluation framework by incorporating additional metrics such as context utilization and answer completeness, providing a more holistic assessment of RAG system performance. These efforts collectively highlight the growing emphasis on rigorous, domain-specific, and multi-dimensional evaluation methods to ensure RAG systems meet the demands of diverse applications while maintaining high retrieval and generation quality standards.
@@ 8.4. Question Answering
RAG techniques combine information retrieval with model-based generation, which is highly suitable for QA systems. In particular, open-domain QA systems usually first require searching for knowledge from the Internet or large-scale databases, then generate the corresponding answers according to the retrieved knowledge. Naturally, given similar questions and corresponding answers as demonstrations which are concatenated into inputs (Wang et al., 2022; Li et al., 2023b; Huang et al., 2023c), generators in RAG can learn the pattern between questions and answers and infer what answers should be. For some specific QA tasks where a set of reference documents is given, retrievers in RAG would retrieve the relevant documents for concatenation, and then generators in RAG would read the context then generate the final answers via the self-attention mechanism (Guu et al., 2020; Lee et al., 2023; Ram et al., 2023b; Asai et al., 2024), which is similar to solving a reading comprehension problem. Besides, Fabbri et al. (Fabbri et al., 2020) focus on designing effective templates for re-organizing the concatenated contexts. Baek et al. (Baek et al., 2023) leverage the knowledge graph to retrieve the related facts for the input questions, then feed their concatenation and inputs into the generators. Instead of directly concatenating texts, another branch of works focuses on joining the retrieval embeddings with input embeddings for the encoder-decoder models (Izacard and Grave, 2021; Sachan et al., 2021; de Jong et al., 2023b; Izacard et al., 2023).
@@
Some works incorporate the external knowledge in the hidden states or the final logits of generators. For the fusion in the hidden states, the key is what kind of knowledge representation would be injected, such as entities (Févry et al., 2020; de Jong et al., 2022), chunks (Borgeaud et al., 2022; Wang et al., 2023f), documents (Cheng et al., 2023a). For the fusion in the logits, most works combine the logits of retrievals and inputs by ensemble techniques (Shi et al., 2023; Guu et al., 2020; Lewis et al., 2020; Mueller et al., 2023).
@@
Instead of designing different knowledge fusions for QA systems, existing works also improve QA systems with RAG from other aspects. Some works (Guo et al., 2019; Paranjape et al., 2022; Lin et al., 2022) use retrieved question-answering pairs as extra training data. Some works optimize the retriever module, e.g., improving the keys’ representation when building the retriever database (Ram et al., 2023a), replacing the indexing with a pre-trained ranking model (Yu et al., 2023b), or enabling retrieving phrases with two queries (Min et al., 2023). Other works focus on accelerating the generation efficiency of RAG. Jong et al. (de Jong et al., 2023a) propose the layer-sparse cross-attention to speed up the decoding. Some works (Asai et al., 2024; Jiang et al., 2023b; Wang et al., 2023c) observe that the retrievals may not always provide useful information during the generation process and learn to determine when to retrieve. Moreover, Sun et al. (Sun et al., 2023) combine the RAG with agents to iteratively reason the final results.
@@ 8.7. Dialogue Systems
Existing works improve the dialogue system with RAG mostly via the query-based fusions. Some works (Li et al., 2021; King and Flanigan, 2023; Cheng et al., 2023b) concatenate the retrieved history conversations with current inputs. Other works (Fan et al., 2021; Liu et al., 2023a; Cheng et al., 2023b) first leverage an encoder to encode the history responses, then feed the concatenated embeddings into a decoder to generate new responses.
@@ 9.1. LLM-based Autonomous Agents
LLM-based autonomous agents are intelligent software systems that leverage the power of LLMs to perform tasks without the need for continuous human intervention (Li et al., 2024; Xi et al., 2023; Wang et al., 2024). These agents use LLMs as a brain or controller (Huang et al., 2024), and extend their abilities through multimodal perception (Xie et al., 2024), tool utilization (Schick et al., 2023) and external memory (Packer et al., 2023). Especially, external long-term memory for agents functions as the knowledge datastore in RAG, which provides agents with the capability to incorporate external knowledge over extended periods. Therefore, applying RAG would be beneficial to access a broader range of information, improving agents’ decision-making and problem-solving abilities (Zhang et al., 2024b). This section explores how LLM-based agents can leverage RAG from two perspectives.
@@ 9.2. Frameworks
Frameworks like Langchain (LangChain, 2023) and LLaMAindex (Liu, 2022) pose a significant impact on enhancing the practical implementation of RAG. Langchain and LLaMAindex exemplify the integration of sophisticated retrieval mechanisms with generative models, facilitating the seamless incorporation of external data into the language generation process. This section will introduce these two representative RAG frameworks in details.
Langchain is a framework designed to augment the capabilities of language models by integrating them with external knowledge sources and databases. It acts as a middleware that facilitates the interaction between language models and various data retrieval systems, enabling a more informed and accurate generation of responses. The core functionality of Langchain involves orchestrating the flow of information from external databases into the generative process of language models, enhancing their ability to leverage context and specific knowledge in their responses. This integration plays a crucial role in enabling language models to perform tasks that require access to up-to-date or detailed information that is not contained within the model’s initial training data.
@@ 10. Discussion and Future Direction
Despite the success of the RAG for natural language processing, some challenges should be considered. This paper highlights these challenges to inspire future research and provides possible future research directions in RAG for NLP.
@@ 10.1. Retrieval Quality
The retrieval quality largely determines the performance of the generator. Park et al. (Park and Lee, 2024) investigated imperfect retrieval’s effect and demonstrated that unanswerability or contradiction of a document set, which frequently leads to hallucinations. However, Ren et al. (Ren et al., 2025) demonstrated that the improvement brought about by the increase in the number of supporting documents is not entirely due to the increase in recall. To this end, using various evaluation methods referred to in Section  7 for retrieval information will be more beneficial to model generation. The retrieval quality involves the following four key factors that must be designed. The first consideration is determining the optimal key to use in the vector database. This process typically involves subjective decision-making and requires human effort to design effectively. The naive idea is to choose inputs for the given tasks, treating each task as a QA problem.
The second is the choice of embedding model. After determining the key, the next step is leveraging embedding models to convert text into vector representations. Models such as BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), or domain-specific embeddings can be crucial to determine how well nuances and contextual meanings are captured. Adapting the embedding model to suit specific data types or queries better can significantly enhance retrieval quality. This requires training the model on domain-specific corpora, including the types of queries and documents the system will encounter.
@@ 10.2. RAG Efficiency
10.3. Discussion for Query-based Fusion in RAG
Two main query-based fusion RAG concerns must be discussed. The first is the performance comparison towards long-context LLMs. Hui et al. (Hui et al., 2024) demonstrated that RAG system and long-context LLMs show similar performance for free-form or knowledge-based tasks (e.g., paper-based and wiki-based Q&A). However, in tasks requiring numerical reasoning (e.g., financial Q&A), long-context LLMs underperform compared to RAG, as the verbose content in long contexts can obscure key facts and hinder precise reasoning. Additionally, RAG tends to outperform long-context mechanisms in smaller LLMs, which struggle to process large volumes of data effectively. The second concern is improving the query-based fusion method, which is prevalent in modern RAG systems. Recent works used prompt engineering strategies (Lewis et al., 2020; Guo et al., 2023), query refinement techniques (Yan et al., 2024), calibration-based strategies (Jiang et al., 2023b; Asai et al., 2024), and iterative RAG (Shao et al., 2023) to improve the query-based fusion RAG. Among these strategies, prompt engineering and calibration offer improved results with minimal overhead, while iterative retrieval-augmented generation (RAG) and query refinement prioritize higher computational costs in exchange for potential accuracy gains. Prompt engineering and query refinement depend on manual tuning, whereas calibration and iterative RAG are more automation-friendly but require high-quality, robust data. Simpler strategies, such as prompting, are easier to deploy but lack flexibility, while more complex methods like iterative RAG provide greater adaptability at the expense of stability. These approaches can be combined (e.g., refined queries with iterative loops and calibrated prompts), though such synergies must carefully balance the compounded drawbacks, such as increased latency and complexity.
@@ 10.5. RAG Training
As introduced in Section 6, RAG training includes two branches of works, RAG with/without datastore update. For RAG without a datastore update, the main challenge is how to jointly optimize all parameters in RAG. This may involve new loss functions with multiple objectives, new optimizations for efficient tuning parameters in the retriever and generator, or other training strategies.
For RAG with datastore update, one challenge is how to align the retrieval representations with the generator’s representations. Although the time cost of the update operation in datastore cannot be ignored, some works (Chen et al., 2022b) reduce the update frequency by asynchronously updating, thus achieving the alignment of knowledge representation and model’s representation. Another challenge is when to retrain/fine-tune the generator in RAG when a new corpus is added. Due to the in-context learning capability of existing LLM-based generators and high training overhead, retraining/fine-tuning the generator or directly inferring the generator becomes a challenging choice for different scenarios. Recently, some efficient training strategies (Hu et al., 2022; Dettmers et al., 2023) have been proposed to accelerate the fine-tuning process, which can be taken into consideration.
@@ 11. Conclusion
In this survey, we delve into the development of RAG within the field of natural language processing. First, this paper introduces the components of RAG and their functionalities. Subsequently, this paper elaborates on each step involved in retriever, discussing the diverse techniques. Furthermore, this paper categorizes the retrieval fusions, evaluating the strengths and weaknesses inherent in each retrieval fusion technique. Besides, this paper discusses the RAG training, including RAG with/without datastore update. Then, this paper presents RAG evaluation and benchmarking, and explores how RAG can be adapted for various NLP tasks and provides practical applications of RAG in real-world scenarios. Conclusively, this paper identifies ongoing challenges and suggests directions for future research to foster advancements in this evolving area.